OpenStudy (wildemo):

Is 2(x-4)^2-19 the vertex form of 2x^2+8x-3?

OpenStudy (welshfella):

expand 2(x - 4)^2 -19 and see what you get

OpenStudy (wildemo):

I see now that it is incorrect. Do you think you could help me with finding the vertex from of 2x^2+8x-3? @welshfella

OpenStudy (welshfella):

ok first divide the first divide first 2 terms in the equation by 2:- 2 [x^2 + 4x ] - 3 Now rewrite the term in the brackets 2[ (x + 2)^2 - 4] - 3 = 2(x + 2)^2 -8 - 3 = 2(x + 2)^2 - 11

OpenStudy (wildemo):

I follow with the first line of number but the second one I'm confused with

OpenStudy (welshfella):

to rewrite (x^2 + 4 ) you use the identity a^2 + bx = (a + b/2)^2 - (b/2)^2

OpenStudy (welshfella):

* rewrite (x^2 + 4x) I should have said

OpenStudy (welshfella):

if you expand (x + 2)^2 you get x^2 + 4x + 4 right? now if we subtarct 4 we get back to x^2 + 4x x^2 + 4x + 4 - 4 = x^2 + 4x its another way of writing x^2 + 4x

OpenStudy (wildemo):

I didn't know at first that if you expand (x + 2)^2 you get x^2 + 4x + 4. but i get it now.

OpenStudy (welshfella):

(x + 2)(x+2) = x(x + 2) + 2(x + 2)

OpenStudy (wildemo):

I get it now wow.

OpenStudy (welshfella):

good